Lazorwolf64:

I don't understand the Angle of Elevation or Angle of Depression, please explain it to me.

Lui0210:

In very very simple terms, when you are on the horizon, the angle of elevation is the angle formed between the horizon and the line ABOVE the horizontal line. The angle of depression is the line that is formed BELOW that horizontal line.

Lui0210:

Lazorwolf64:

Thank you for the visual picture! I'm more confused about how would you solve it? Let's say You're sitting on top of a building and you look down and see your friend on the street. If the angle of depression to your friend is 57 degrees, and the building is 200 feet tall, how far away from the bottom of the building is your friend?

Lui0210:

It's just giving directions on how the angle(s) should be drawn in order to get the right calculations. So looking down would be angle of depression AT the horizon, the horizon being the first person's POV. And 200 ft being one of the sides of the triangle. So it would look something like this.

Lui0210:

We're going to assume the building is straight, so the height of the building and the angle of depression add up to make a right angle.

Lazorwolf64:

@lui0210 wrote:
We're going to assume the building is straight, so the height of the building and the angle of depression add up to make a right angle.
So then would I use Tangent? (Tan(57) =200/x? how would I solve that?

Lui0210:

This is why depression vs elevation is important. You would not use the angle of depression, that is just a reference. You would actually use the angle that is between the building and The Angle of Depression. So, x = 90-57 Then use the tan function like you did earlier

Lazorwolf64:

@lui0210 wrote:
This is why depression vs elevation is important. You would not use the angle of depression, that is just a reference. You would actually use the angle that is between the building and The Angle of Depression. So, x = 90-57 Then use the tan function like you did earlier
I'm so sorry, could you explain that further?

toga:

The angle of elevation is like when you look up at something, and it's the angle between your line of sight and the horizontal line. This angle helps figure out how high something is above a horizontal line or plane. And the angle of depression is like when you look down at something, and it's the angle between your line of sight and the horizontal line. This angle helps figure out how deep or high something is below a horizontal line or plane. Hope that helps! tag me if you have any other questions.